# Stochastic Kinetics of Transcription: Analysis and Computation

### v1

@@ 3.2 Special Case I: Telegraph model
Taking $D_{0}=\begin{pmatrix}-k_{1}-\lambda&k_{1}\\ k_{2}&-k_{2}\end{pmatrix}$ and $D_{1}=\text{diag}[\lambda,0]$, our model (1) reduces to the classic Telegraph model [4], the pioneering work in the field of stochastic gene expression. It is clear that $\bm{\pi}=\left[k_{2}/(k_{1}+k_{2}),k_{1}/(k_{1}+k_{2})\right]^{T}$ and
@@
where  $\mathllap{{}_{1}}F_{1}(a,b;z):=\sum_{n=0}^{\infty}\frac{a^{(n)}}{b^{(n)}}\frac{z^{n}}{n!}$ denotes the confluent hypergeometric function. (12) is consistent with previous result [4, 8]. A detailed calculation can be found in the Supplementary Material [21].
Although formulation of $\{\mathcal{P}_{n}\}_{n\in\mathbb{N}}$ in (12) seems formally compact, the confluent hypergeometric function is, in essence, an abbreviation of an infinite series. Evaluating the value of confluent hypergeometric function at given point possesses approximately the same computational complexity as one calculation in (7), and it involves truncation error as well. From this standpoint, (12) holds limited advantage, both mathematically and computationally, over reconstructing $\{\mathcal{P}_{n}\}_{n\in\mathbb{N}}$ from binomial moments according to (7). Likewise, this observation holds for other known results involving generalized hypergeometric function [8, 9, 10, 17].
@@ 3.3 Special Case II: Under Renewal Condition
$\text{if}\;\;\sum_{k=1}^{N}b_{i,k}\neq 0,\;\;\text{then}\;\;\frac{b_{i,k}}{\sum_{k=1}^{N}b_{i,k}}\;\;\text{is independent of}\;\;i.$ (13)
Assuming $D_{1}\neq\bm{0}_{N\times N}$, the renewal condition (13) is equivalent to $\text{rank}\,D_{1}=1$. Under the renewal condition, our model (1) can be interpreted otherwise as a queueing system $GI/M/\infty$ with probability density function of inter-arrival time [17]
@@
where $\bm{\kappa}=(\kappa_{j})^{T}=(\frac{b_{i,j}}{\sum_{j=1}^{N}b_{i,j}})^{T}$ as long as $\sum_{j=1}^{N}b_{i,j}\neq 0$.
@@
[…]
Assuming nonsingularity of $D_{0}$, binomial moments given in (15) can be reduced to our result given in (5). The details can be found in the Supplementary Material [21]. We note that, under the assumption $D$ is irreducible and $\text{rank}\,D_{1}=1$, $D_{0}$ is nonsingular, which is equivalent to convergence of $\alpha$. A proof is given in the Supplementary Material [21], which can be seen as a generalization of Taussky theorem [23].
@@ 3.4 Computational Performance
Figure 3: Computation Time of Binomial Moments and SSA: Computation time is evaluated for increasing model complexity and different approaches. As dimension of $D_{0}$ or $D_{1}$ increases, number of reaction pathways escalates, making the model more complicated. In the above illustration, computation time pertains to an exemplar where $a_{i,j}=0(i\neq j)$ and $b_{i,j}=i$. In the left figure, three different approaches are employed to models with dimension $i\,(3\leq i\leq 8)$; in the right one, two approaches are employed to model with dimension $5j\,(1\leq j\leq 10)$. Curve of our result is computation time of the first $200$ binomial moments based on (5). Curve of renewal process is computation time of the first $200$ binomial moments based on (15). SSA is performed for $1000$ trajectories and truncated at dimensionless time $t=100$ (GillesPy2 [27] is implemented with Python solver). Time is evaluated for three replicate computations and averaged to mitigate randomness.
@@
Our result for (1) is performed mainly in the framework of matrix computation, which is one of the most efficient and well-developed area in scientific computation. As we illustrate in Figure 3, our result stands out among several existing approaches in computational efficiency, where we compute the first $200$ binomial moments based on (5) and (15), respectively, and implement SSA. Separation of time scales can be distinctly observed with respect to our result, the renewal process approach and SSA. In particular, our result remains computationally feasible for high-dimensional model while other approaches certainly fail. Moreover, our result also admits explicit error estimation, which is, as far as we know, unattainable through other existing methods. Additionally, accuracy, stability and further improvement on efficiency can be pursued by exploiting the structure of matrix in (5), especially $k\delta\bm{I}_{N}-D$. For example, $k\delta\bm{I}_{N}-D$ is diagonally dominant, thus can perform LU factorization without pivoting (Theorem $4.1.1$ in [26]). It involves $\frac{8}{3}N^{3}$ flops to calculate $\left(k\delta\bm{I}_{N}-D\right)^{-1}$ for given $k$ and $\frac{20}{3}(m-1)N^{3}$ flops to calculate $B_{m}$ in general. Recursive structure in (5) can also be utilized to reduce overall complexity.
In comparison, SSA has two notable shortcomings. Firstly, SSA is computationally intensive and thus impractical to implement for high-dimensional reaction system. The second is that SSA relies heavily on empirical insights. Number of simulations and truncation time point to generate a stationary distribution are both empirically determined, hence the implementation requires repeated trials. The renewal process approach also has several shortcomings. Primarily, Laplace transform of inter-arrival time probability density distribution can be mathematically challenging to obtain given specific gene expression model. Inter-arrival time distribution also conceals physical interpretation of gene-state architecture while parameters in (1) have clear biophysical meanings. Additionally, results based on renewal process is computational infeasible excluding a few special cases. In general, computational complexity become exceedingly high if high-order binomial moment is involved.
4 Mathematical Description of Noise in Gene Expression
@@
Noise in gene expression within a supposedly identical ensemble arises from multiple sources [2, 3]. Quantitative understanding of generation and propagation of noise is one of the primary objectives of studying stochastic gene expression, and noise in transcription can predominantly contribute to protein-level fluctuation [29]. Experimentally, noise can be categorized into intrinsic and extrinsic noise. Intrinsic noise is the fluctuation originating from inherent stochasticity of biochemical reactions. Deterministic description fails since number of molecules involved in gene expression can be relatively low. Our model (1) provides a general framework to investigate origins and consequences of intrinsic noise in gene expression. Based on our result (5), Fano factor, the ratio of variance to mean of mRNA copy number, is given by
$FF=\frac{2B_{2}+B_{1}-B_{1}^{2}}{B_{1}}=\frac{\bm{\pi}^{T}D_{1}(\delta\bm{I}_{N}-D)^{-1}D_{1}\bm{1}}{\bm{\pi}^{T}D_{1}\bm{1}}+1-\frac{\bm{\pi}^{T}D_{1}\bm{1}}{\delta}.$ (16)
@@
$\eta^{2}=\frac{2B_{2}+B_{1}-B_{1}^{2}}{B_{1}^{2}}=\frac{\delta\bm{\pi}^{T}D_{1}(\delta\bm{I}_{N}-D)^{-1}D_{1}\bm{1}}{(\bm{\pi}^{T}D_{1}\bm{1})^{2}}+\frac{\delta}{\bm{\pi}^{T}D_{1}\bm{1}}-1.$ (17)
For specific gene-state structure where $(\delta\bm{I}_{N}-D)^{-1}$ has explicit expression, further reduction of Fano factor and intrinsic noise may be performed. Consistency of (16) and (17) with previous results can be verified [9, 17].
@@
Extrinsic noise refers to fluctuation caused by environmental differences among cells, and can be further classified as local and global extrinsic noise [2]. Although intrinsic noise can easily be grounded in mathematical formulation, as our model (1) does, incorporating extrinsic noise into chemical mater equation is conventionally considered to be formidable. One insightful approach is to utilize the parametric nature of the model [19]. Parameters, which are $D_{0},D_{1}$ and $\delta$ in our case, are assumed to follow some probability distribution with density function $\rho(\bm{\theta})$, representing effects of extrinsic fluctuation. Since binomial moments have concise expression given by (5), posing the perturbation on binomial moments and then reconstruct probability distribution is more tractable than posing perturbation directly on probability distribution as in [19]. Thereby, the expected binomial moments of mRNA molecules subject to extrinsic noise can be formulated as
[…]
@@
where $\bm{\theta}$ abbreviates the vector of $12N^{2}-N+$ 1 independent parameters in our model (1) and $\Omega$ denotes the parameter space $[0,\infty]^{2N^{2}-N+1}$. Substituting (18) into (7), perturbed probability distribution can be reconstructed. For specific $\rho(\bm{\theta})$ and simplified model, (18) may be reduced to closed form.
@@
This framework can be used to apprehend the joint effects of intrinsic and extrinsic noise, which has been performed for some special cases of our model (1) [8, 19]. Heavy-Tailed Law [19] asserts that intrinsic noise alone cannot generate mRNA distributions that is heavy tailed, namely, the moment generating function [$\mathcal{M}\left[\{\mathcal{P}_{n}\}_{n\in\mathbb{N}}\right](t)$ becomes infinity for $t>0$. Heavy-Tailed Law has been proposed intuitively for Leaky Telegraph model [19] and a generalized model [8]. In this Letter, we provide a mathematically rigorous proof of Heavy-Tailed Law in our model (1), generalizing previous results. More preciously, we prove the following inequality based on (8).
@@
For Poisson distribution, moment generating function is finite for any $t>0$, therefore so is $\{\mathcal{P}_{n}\}_{n\in\mathbb{N}}$ according to (19). Therefore $\{\mathcal{P}_{n}\}_{n\in\mathbb{N}}$ is never heavy tailed. It follows that experimentally observed heavy-tailed distribution of mRNA copy number implies collaborative effect of intrinsic and extrinsic noise [19]. The mathematical proof of (19) is in Supplementary Material [21].
@@ 5 Discussion
Stochastic gene expression is a central problem in biophysics. Numerous mathematical theories have been developed to provide quantitative insights into specific models. In this Letter, we study a general kinematic model of transcription. The exact time-dependent solution to the chemical master equation is constructed. A computationally feasible approach based on binomial moments is presented and thoroughly analyzed. We examine two special cases and perform SSA to verify our results. Collaborative effect of intrinsic and extrinsic noise are discussed. A mathematically rigorous proof of generalized Heavy-Tailed Law is also provided. Compared with other existing methods, our approach features generality, exceptional computational efficiency as well as explicit error estimation. Results presented in this Letter boost further analysis of origin and variation of noise. Dependence of noise and Fano factor on different parameters should be further investigated. Note also that our model is theoretically motivated. When modeling specific type of gene, physical background should be exploited to constrain model complexity appropriately. Statistical inference can further be employed by fitting our model to experimental data, and finally reveal the underlying mechanisms of transcription.
@@ Acknowledgments
Y. L. is supported by FDUROP (Fudan Undergraduate Research Opportunities Program). Y. Z. is supported by National Key R&\&&D Program of China (2024YFA1012401), the Science and Technology Commission of Shanghai Municipality (23JC1400501), and Natural Science Foundation of China (12241103).

### v2

@@ 3.1 Special Case I: Telegraph model
With $D_{0}=\begin{pmatrix}-k_{1}-\lambda&k_{1}\\ k_{2}&-k_{2}\end{pmatrix}$ and $D_{1}=\text{diag}[\lambda,0]$, our model (1) reduces to the classic Telegraph model, the pioneering mathematical formulation in the field of stochastic gene expression. It is clear that $\bm{\pi}=\left[k_{2}/(k_{1}+k_{2}),k_{1}/(k_{1}+k_{2})\right]^{T}$ and
@@
where  $\mathllap{{}_{1}}F_{1}(a,b;z):=\sum_{n=0}^{\infty}\frac{a^{(n)}}{b^{(n)}}\frac{z^{n}}{n!}$ denotes the confluent hypergeometric function. (12) is consistent with previous result [13, 30]. A detailed calculation can be found in Appendix F.
Although formulation of $\{\mathcal{P}_{n}\}_{n\in\mathbb{N}}$ in (12) seems formally compact, the confluent hypergeometric function is, in essence, an abbreviation of an infinite series. Evaluating the value of confluent hypergeometric function at given point possesses approximately the same computational complexity as one calculation in (7), and it involves truncation error as well. From this standpoint, (12) holds limited advantages, both mathematically and computationally, over reconstructing $\{\mathcal{P}_{n}\}_{n\in\mathbb{N}}$ from binomial moments according to (7). Likewise, this observation holds for other known results involving generalized hypergeometric function [4, 13, 16, 37].
@@ 3.2 Special Case II: Under Renewal Condition
$\text{When}\;\;\sum_{k=1}^{N}b_{i,k}\neq 0,\;\;\frac{b_{i,k}}{\sum_{k=1}^{N}b_{i,k}}\;\;\text{is independent of}\;\;i.$ (13)
Assuming $D_{1}\neq\bm{0}_{N\times N}$, renewal condition (13) is equivalent to $\text{rank}\,D_{1}=1$. Under renewal condition, our model (1) can be interpreted otherwise as a queueing system $GI/M/\infty$ with probability density function of inter-arrival time [37]
@@
where $\bm{\kappa}=(\kappa_{j})^{T}:=(\frac{b_{i,j}}{\sum_{j=1}^{N}b_{i,j}})^{T}$ as long as $\sum_{j=1}^{N}b_{i,j}\neq 0$.
@@
where $\mathcal{L}[f](s):=\int_{0}^{\infty}\mathrm{e}^{-st}f(t)\mathrm{d}t$ is the Laplace transform of $f(t)$ and $\alpha:=\int_{0}^{\infty}tf(t)\mathrm{d}t$ is the expectation of inter-arrival time.
Assuming nonsingularity of $D_{0}$, binomial moments given in (15) can be reduced to our result given in (5). The details can be found in Appendix G. We note that, under the assumption $D$ is irreducible and $\text{rank}\,D_{1}=1$, $D_{0}$ is nonsingular, which is equivalent to convergence of $\alpha$. A proof is presented in Appendix H, which can be seen as an analogue to Taussky theorem [15].
@@ 3.3 Computational Performance
Figure 3: Computation Time of Binomial Moments and SSA: Computation time is evaluated for increasing model complexity and different approaches. As dimension of $D_{0}$ or $D_{1}$ increases, number of reaction pathways escalates, making the model more complicated. In the above illustrations, computation time pertains to exemplars where $a_{i,j}=0(i\neq j)$ and $b_{i,j}=i$. In the left panel, three different approaches are employed to models with dimension $k\,(3\leq k\leq 8)$; in the right one, two approaches are employed to models with dimension $5l\,(1\leq l\leq 10)$. Curve of our result is computation time of the first $200$ binomial moments based on (5). Curve of renewal process is computation time of the first $200$ binomial moments based on (15). SSA is performed for $1000$ trajectories and truncated at dimensionless time $t=100$ (GillesPy2 [24] is implemented with Python solver). Time is evaluated for three replicate computations and averaged to mitigate randomness.
@@
Our result (5) is performed mainly in the framework of matrix computation, which is one of the most efficient and well-developed area in scientific computation. As we illustrate in Figure 3, our result stands out in computational efficiency among several existing approaches, where we compute the first $200$ binomial moments based on (5) and (15), respectively, and implement SSA. Separation of time scales can be distinctly observed with respect to our result, the renewal process approach and SSA. In particular, our result remains computationally feasible for high-dimensional model while other approaches may fail. Moreover, our result also admits explicit error estimation (9), which is, as far as we know, unattainable through other existing methods. Additionally, accuracy, stability and further improvement on efficiency can be pursued by exploiting the structure of matrix in (5), especially $k\delta\bm{I}_{N}-D$. For example, $k\delta\bm{I}_{N}-D$ is diagonally dominant, thus can perform LU factorization without pivoting (Theorem $4.1.1$ in [8]). It involves $\frac{8}{3}N^{3}$ flops to calculate $\left(k\delta\bm{I}_{N}-D\right)^{-1}$ for given $k$ and $\frac{20}{3}(m-1)N^{3}$ flops to calculate $B_{m}$ in general. Recursive structure in (5) can also be utilized to reduce overall complexity. In comparison, SSA has two notable shortcomings. Firstly, SSA is computationally intensive and thus impractical to implement for high-dimensional reaction system. The second is that SSA relies heavily on empirical insights. To generate a stationary distribution, number of simulations and truncation time point are both empirically determined, hence the implementation requires repeated trials. The renewal process approach also has shortcomings. Primarily, Laplace transform of inter-arrival time probability density function can be mathematically challenging to obtain given specific gene expression model. Inter-arrival time probability density function also conceals physical interpretation of gene-state architecture while parameters in (1) have clear physical meanings. Additionally, results based on renewal process are computational infeasible excluding a few special cases, because, in general, computational complexity becomes exceedingly high if high-order binomial moments are involved.
4 Fluctuation of mRNA Copy Number in Transcription
@@
Noise in gene expression within a supposedly identical ensemble arises from multiple sources [26, 32]. Quantitative understanding of generation and propagation of noise is one of the primary objectives of modeling stochastic gene expression, and noise in transcription can predominantly contribute to protein-level fluctuation [28]. Experimentally, noise can be categorized into intrinsic and extrinsic noise. Intrinsic noise is the fluctuation originating from inherent stochasticity of biochemical reactions. Deterministic description fails since number of molecules involved in gene expression can be relatively low. Our model (1) provides a general framework of investigating origins and consequences of intrinsic noise in transcription. Based on our result (5), Fano factor, the ratio of variance to mean of mRNA copy number, is given by
$FF:=\frac{2B_{2}+B_{1}-B_{1}^{2}}{B_{1}}=\frac{\bm{\pi}^{T}D_{1}(\delta\bm{I}_{N}-D)^{-1}D_{1}\bm{1}}{\bm{\pi}^{T}D_{1}\bm{1}}+1-\frac{\bm{\pi}^{T}D_{1}\bm{1}}{\delta}.$ (16)
@@
$\eta^{2}:=\frac{2B_{2}+B_{1}-B_{1}^{2}}{B_{1}^{2}}=\frac{\delta\bm{\pi}^{T}D_{1}(\delta\bm{I}_{N}-D)^{-1}D_{1}\bm{1}}{(\bm{\pi}^{T}D_{1}\bm{1})^{2}}+\frac{\delta}{\bm{\pi}^{T}D_{1}\bm{1}}-1.$ (17)
For specific gene-state structure where $(\delta\bm{I}_{N}-D)^{-1}$ has explicit expression, further reduction of Fano factor and intrinsic noise may be performed. Consistency of (16) and (17) with previous results can be verified [4, 37]. Derivation of (16) and (17) is in Appendix I. As illustrated in Figure 4, intrinsic noise converges to zero as model size increases, implying that kinetics of mRNA molecules tends towards deterministic description in these exemplars. Additionally, Fano factor is greater than $1$ except for the trivial $n=1$ case according to Figure 4, meaning that distribution of mRNA copy number is super-Poissonian in these exemplars.
Figure 4: Variation of Intrinsic Noise and Fano Factor: Intrinsic noise and Fano factor are computationally evaluated for models of order $n$($1\leq n\leq 100$), with parameters $a_{i,j}=0(i\neq j)$ and $b_{i,j}=i$. In the left panel, intrinsic noise $\eta^{2}$ is computed based on (17) and $-\ln(\eta^{2})$ is taken for plotting. In the right panel, Fano factor is computed based on (16).
@@
Extrinsic noise refers to fluctuation caused by environmental differences among cells, and can be further classified as local and global extrinsic noise [32]. Although intrinsic noise can easily be grounded in mathematical formulation, as our model (1) does, incorporating extrinsic noise into chemical mater equation is conventionally considered to be formidable. One insightful approach is to utilize the parametric nature of the model [10]. Parameters, which are $D_{0},D_{1}$ and $\delta$ in our case, are assumed to follow given probability distribution with density function $\rho(\bm{\theta})$, representing effects of extrinsic perturbation. Since binomial moments have concise expression given by (5), posing the perturbation on binomial moments and then reconstruct probability distribution is more tractable than posing perturbation directly on probability density function as in Ref. [10]. Thereby, the expected binomial moments of mRNA molecules subject to extrinsic noise can be formulated as
$\langle B_{m}\rangle_{\rho}=\int_{\Omega}B_{m}(\bm{\theta})\rho(\bm{\theta})\mathrm{d}\bm{\theta},$ (18)
@@
where $\bm{\theta}$ abbreviates the vector of $12N^{2}-N+$ 1 independent parameters in our model (1) and $\Omega$ denotes the parameter space $[0,\infty]^{2N^{2}-N+1}$. Substituting (18) into (7), perturbed probability density function can be reconstructed. For specific $\rho(\bm{\theta})$ and simplified model, (18) may be reduced to closed form. In general, (18) can also be numerically evaluated by Monte Carlo integration since computation of $B_{m}$ is relatively fast.
@@
This framework can be used to apprehend the joint effects of intrinsic and extrinsic noise, which has been performed for some special cases of our model (1) [10, 13]. Heavy-Tailed Law asserts that intrinsic noise alone cannot generate mRNA distributions that is heavy tailed, namely, the moment generating function [$\mathcal{M}\left[\{\mathcal{P}_{n}\}_{n\in\mathbb{N}}\right](t)$ becomes infinity for any $t>0$. Heavy-Tailed Law has been proposed intuitively for Leaky Telegraph model [10] and a generalized model [13]. Here we provide a mathematically rigorous proof of Heavy-Tailed Law in our model (1), generalizing previous results. Denote Poisson distribution with parameter $\Lambda$ by $\text{Pois}(\Lambda)$. We prove the following inequality based on (8).
@@
For Poisson distribution, moment generating function is finite for any $t>0$, therefore so is $\{\mathcal{P}_{n}\}_{n\in\mathbb{N}}$ according to (19). It follows that experimentally observed heavy-tailed distribution of mRNA copy number implies collaborative effect of intrinsic and extrinsic noise [10]. The mathematical proof of (19) is in Appendix J.
@@ 5 Discussion
Stochastic gene expression is a central problem in biophysics. Numerous mathematical theories have been developed to provide quantitative insights into specific models. This article establishes a general framework for studying stochastic transcription kinetics and related problems. While pursuing the utmost theoretical generality, we construct the exact time-dependent solution to CME and present a computationally feasible approach based on binomial moments. We unify existing studies on transcription kinetics and discuss the combined effects of intrinsic and extrinsic noise. Additionally, a mathematically rigorous proof of generalized Heavy-Tailed Law is provided. Compared with other existing methods, our approach features generality, exceptional computational efficiency as well as explicit truncation error estimation. Results presented in this article boost further analysis of origin and variation of noise. Dependence of noise and Fano factor on different parameters should be further investigated. Note also that our model is theoretically motivated. When modeling specific type of gene, physical background should be exploited to constrain model complexity appropriately. Statistical inference can further be employed by fitting our model to experimental data, eventually revealing the underlying mechanisms of transcription.
@@ Acknowledgments
Y. L. is supported by FDUROP (Fudan Undergraduate Research Opportunities Program)(24260). Y. Z. is supported by National Key R&\&&D Program of China (2024YFA1012401), the Science and Technology Commission of Shanghai Municipality (23JC1400501), and Natural Science Foundation of China (12241103).
Author Declaration
Conflict of Interest
The authors have no conflicts to disclose.
Author Contributions
Yuntao Lu: Conceptualization (equal); Formal analysis (lead); Software (lead); Visualization (lead); Writing - original draft (lead); Writing - review & editing (equal). Yunxin Zhang: Conceptualization (equal); Project administration (lead); Supervision (lead); Writing - review & editing (equal).
Data Availability Statement
The data that supports the findings of this study are available within the article.